# A Volumetric Approach to Privacy of Dynamical Systems

### v3

@@ 3.2 Inference Attack Approximation
The tightest recursive inference interval from (2.3) to (13) can be computed via
$\mathcal{M}_{k-1|k}^{x}=$ $\Psi\left(A_{1}^{-1}\right)\mathcal{M}^{x}_{k|k}+\Psi\left(-A_{1}^{-1}A_{2}\right)\mathcal{Y}_{k-1|k-1}$
$+\Psi\left(-A_{1}^{-1}B_{1}\right)\mathcal{W}_{k|k}^{x},$ (20)
$\mathcal{M}_{k-1|k}^{y}=$ $\Psi\left(A_{2}^{-1}\right)\mathcal{M}^{x}_{k|k}+\Psi\left(-A_{2}^{-1}A_{1}\right)\mathcal{Y}_{k-1|k-1}$
$+\Psi\left(-A_{2}^{-1}B_{1}\right)\mathcal{W}_{k|k}^{x},$ (21)
@@
$\!\!\!\!\!\!\mathcal{M}_{k|k-1}^{x}\!=\!\Psi\!\left(A_{1}\right)\!\mathcal{X}_{k-1|k}\!+\!\Psi\!\left(A_{2}\right)\!\mathcal{Y}_{k-1|k}\!+\!\Psi\!\left(B_{1}\right)\!\mathcal{W}_{k}^{x},$ (28)
@@
$\!\!\!\!\!\!\!\!\mathcal{Y}_{k|k}=\Psi\left(A_{3}\right)\mathcal{X}_{k-1|k}+\Psi\left(A_{4}\right)\mathcal{Y}_{k-1|k}+\Psi\left(B_{2}\right)\mathcal{W}_{k}^{y},$ (32)
@@
Also, the tightest prior inference set of $Y_{k}$ is
$\!\!\!\!\!\!\!\mathcal{Y}_{k|k-1}\!=\!\Psi\left(\!A_{3}\!\right)\!\mathcal{X}_{k-1|k-1}\!+\!\Psi\left(\!A_{4}\!\right)\!\mathcal{Y}_{k-1|k-1}\!+\!\Psi\left(\!B_{2}\!\right)\!\mathcal{W}_{k}^{y},$ (33)
@@
According to Lemma 2, the inference attack can be approximately computed using intervals. It can be verified that the intervals computed in Lemma 2 are tight when $X_{k}$ and $Y_{k}$ are scalar variables.
3.3 Properties of the Inference Attack
The inference attack exhibits certain properties. For instance, the radius of the uncertainty $\mathcal{Y}_{k|k}$, i.e., $p_{k|k}^{y}=\overline{Y}_{k|k}-\underline{Y}_{k|k}$ is bounded by a function of the radius of the process noise and the observation set, as stated below.
@@ 3.3 Properties of the Inference Attack
where $\overline{p}^{x}\geqslant p^{m,x}_{j|j}$ for any $j\geqslant 0$, $p^{m,x}_{k|k}$, $p^{w,x}_{k}$ and $p^{w,y}_{k}$ are radii of $\mathcal{M}_{k|k}^{x}$, $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$, respectively.
Proof 3.2.
@@
Since the volume of $\mathcal{Y}_{k|k}$ is the sum of $p_{k|k}^{y}$, the volume $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ is also bounded by a function of $\overline{p}^{x}$. As a result, if the value of elements in $\overline{p}^{x}$ are small, i.e., if the observation set $\mathcal{M}_{k|k}^{x}$ is small, then the privacy level is low, and the adversary experiences less uncertainty after performing the inference attack.
@@
$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}\diagdown\mathcal{Y}_{k|k}.$
Proposition 4.
If $\mathcal{Y}_{k|k}$ is a subset of $\mathcal{Y}_{k|k-1}$, then the volume of $\Delta\mathcal{Y}_{k|k}$ is
$\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)=\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)\!.$
Proof 3.3.
Since $\mathcal{Y}_{k|k}$ is the subset of $\mathcal{Y}_{k|k-1}$, the amount of uncertainty reduction $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is equal to $\left\|\mathcal{Y}_{k|k-1}-\mathcal{Y}_{k|k}\right\|_{1}$, which can be computed as follows,
$\left\|\mathcal{Y}_{k|k-1}-\mathcal{Y}_{k|k}\right\|_{1}$
$=$ $\sum_{i=1}^{n_{y}}\left(\overline{Y}_{k|k-1}\left(i\right)-\overline{Y}_{k|k}\left(i\right)+\underline{Y}_{k|k}\left(i\right)-\underline{Y}_{k|k-1}\left(i\right)\right)$
$=$ $\sum_{i=1}^{n_{y}}{\left(\overline{Y}_{k|k-1}\left(i\right)-\underline{Y}_{k|k-1}\left(i\right)\right)}-\sum_{i=1}^{n_{y}}{\left(\overline{Y}_{k|k}\left(i\right)-\underline{Y}_{k|k}\left(i\right)\right)}$
$=$ $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right),$
where $\left(\overline{Y}_{k|k-1}\left(i\right)-\overline{Y}_{k|k}\left(i\right)+\underline{Y}_{k|k}\left(i\right)-\underline{Y}_{k|k-1}\left(i\right)\right)$ is the length of i-th interval of $\Delta\mathcal{Y}_{k|k}$.
According to Proposition 4, to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$, it is equivalent to reduce the amount of uncertainty reduction $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ since the amount of prior uncertainty $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$ is fixed at time $k$. As shown in the next theorem, the amount of uncertainty reduction, i.e., $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$, is bounded by the new information extracted from $\mathcal{M}^{x}_{k|k}$.
@@
$\!\!\!\!\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)\!=\!\left\|\Psi\!\left(A_{3}\right)\!\Delta\mathcal{X}_{k-1|k}+\Psi\!\left(A_{4}\right)\!\Delta\mathcal{Y}_{k-1|k}\right\|_{1}\!,\!\!$ (38)
@@
$\!\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)\!\!\leqslant\!\left\|\!A_{3}\!\right\|\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!+\!\left\|\!A_{4}\!\right\|\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)\!,$
where $\Delta\mathcal{X}_{k-1|k}=\mathcal{X}_{k-1|k-1}-\mathcal{X}_{k-1|k}$ and $\Delta\mathcal{Y}_{k-1|k}=\mathcal{Y}_{k-1|k-1}-\mathcal{Y}_{k-1|k}$ are the adversary’s uncertainty reduction of $X_{k-1}$ and $Y_{k-1}$, and $\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}$ quantifies the difference in central estimation with and without considering $\mathcal{X}_{k|k}$, $\left\|A\right\|\frac{\Delta}{=}\sum_{i,j}^{n,m}{\left|a_{i,j}\right|}$ is the absolute value of $A$.
Proof 3.4.
@@
According to Theorem 5, the reduction of uncertainty $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is highly correlated with the amount of information that the adversary extracts from the observation set $\mathcal{M}_{k|k}^{x}$. Furthermore, with Proposition 4 and Theorem 5, we have the following lemma to bound the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k-1|k}\right)$.
@@
$\mathrm{Vol}\!\left(\!\mathcal{Y}_{k|k-1}\!\right)\!-\!\left\|\!A_{3}\!\right\|\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!-\!\left\|\!A_{4}\!\right\|\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)$
@@
Therefore, on the one hand, one can reduce the extracted information $\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)$ and $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right)$ to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$. On the other hand, when the private level is high, we also avoid the adversary to update its central estimation $c_{k|k}^{y}$ due to the small value of $\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}$.
@@ 4 Privacy Filter Design Problem Using the Volumetric Privacy measure
In this section, we address the privacy filter design problem, where the filter outputs an appropriate observation set that balances the trade-off between the data utility of the public state and the privacy protection of the private state.
@@ 4.1 The Structure of Privacy Filter
We begin by defining the decision domain of the privacy filter as follows. At time $k$, given the last decision set $\mathcal{X}_{k-1|k-1}$ and the private set $\mathcal{Y}_{k-1|k-1}$, the tightest interval of $X_{k}$ can be computed via Lemma 9, i.e.,
$\!\!\mathcal{X}_{k|k-1}\!=\!\Psi\!\left(\!A_{1}\!\right)\!\mathcal{X}_{k-1|k-1}\!+\!\Psi\!\left(\!A_{2}\!\right)\!\mathcal{Y}_{k-1|k-1}\!+\!\Psi\!\left(B_{1}\right)\!\mathcal{W}_{k}^{x}\!,$
which contains all possible public states that can be reached from any states in $\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k-1}$. Therefore, $\mathcal{X}_{k|k-1}$ is the maximum observation set $\mathcal{M}^{x}_{k|k}$ that the filter can release, i.e., $\mathcal{M}^{x}_{k|k}\subset\mathcal{X}_{k|k-1}$. To maintain the high data utility, the observation set has to satisfy the following constraint,
@@
We next consider the privacy filter drawn in Fig.2, which first randomly generates a set $\mathcal{S}_{k|k}^{x}\subseteq\mathcal{M}^{x}_{k|k}$ that contains the actual public state $X_{k}=x_{k}$, and then optimizes the observation set $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$ based on linear programming. We will demonstrate that the adversary cannot reduce its uncertainty of the public state by reversing the proposed filtering policy.
Figure 2: The structure of privacy filter.

### v4

@@ 3.2 Inference Attack Approximation via Interval Analysis
The recursive inference interval from (6) to (9) can be computed via
$\mathcal{M}_{k-1|k}^{x}$ $=\Psi\left(A_{1}^{-1}\right)\mathcal{M}^{x}_{k|k}\oplus\Psi\left(-A_{1}^{-1}A_{2}\right)\mathcal{Y}_{k-1|k-1}$
$\oplus\Psi\left(-A_{1}^{-1}B_{1}\right)\mathcal{W}_{k|k}^{x},$ (24)
$\mathcal{M}_{k-1|k}^{y}$ $=\Psi\left(A_{2}^{-1}\right)\mathcal{M}^{x}_{k|k}\oplus\Psi\left(-A_{2}^{-1}A_{1}\right)\mathcal{X}_{k-1|k-1}$
$\oplus\Psi\left(-A_{2}^{-1}B_{1}\right)\mathcal{W}_{k|k}^{x},$ (25)
@@
$\!\!\!\!\!\!\mathcal{M}_{k|k-1}^{x}\!=\!\Psi\!$ $\left(A_{1}\right)\!\mathcal{X}_{k-1|k}\!\oplus\!\Psi\!\left(A_{2}\right)\!\mathcal{Y}_{k-1|k}\!\oplus\!\Psi\!\left(B_{1}\right)\!\mathcal{W}_{k}^{x},$ (32)
@@
$\!\!\!\!\!\!\!\!\mathcal{Y}_{k|k}\!=\!\Psi\!$ $\left(A_{3}\right)\!\mathcal{X}_{k-1|k}\!\oplus\!\Psi\!\left(A_{4}\right)\mathcal{Y}_{k-1|k}\!\oplus\!\Psi\!\left(B_{2}\right)\mathcal{W}_{k}^{y}\!,\!$ (36)
@@
Also, the prior inference set of $Y_{k}$ is
$\!\!\!\!\!\mathcal{Y}_{k|k\!-\!1}\!=\!\Psi\!\left(\!A_{3}\!\right)\!\mathcal{X}_{k-1|k-1}\!\oplus\!\Psi\!\left(\!A_{4}\!\right)\!\mathcal{Y}_{k-1|k-1}\!\oplus\!\Psi\!\left(\!B_{2}\!\right)\!\mathcal{W}_{k}^{y}\!,\!\!$ (37)
@@
Given the interval-based inference approach described in Lemma 6, the computational complexity of the inference attack can be characterized as follows.
Proposition 7.
The computational complexity of the inference attack via interval analysis is $\mathcal{O}(n^{3})$.
Proof 3.3.
The dominant operation in the inference attack involves matrix multiplication. Since the matrix $\Psi(\star)$ has dimensions $(2n\times 2n)$, the corresponding computational complexity is $\mathcal{O}(n^{3})$.
Although the matrix multiplication with large $n$ can still be computationally demanding, the complexity of interval-based inference is substantially lower and remains constant over time, in sharp contrast to the exponentially growing complexity of CCG-based inference.
3.3 Properties of the Interval Inference Attack
The inference attack exhibits certain properties. For instance, the radius of the uncertainty $\mathcal{Y}_{k|k}$, i.e., $p_{k|k}^{y}=\overline{Y}_{k|k}-\underline{Y}_{k|k}$ is bounded by a function of the radius of the disturbance and the observation set, as stated below.
@@ 3.3 Properties of the Interval Inference Attack
where $\overline{p}^{x}\geqslant p^{m,x}_{j|j}$ for any $j\geqslant 0$, $p^{m,x}_{k|k}$, $p^{w,x}_{k}$ and $p^{w,y}_{k}$ are radii of $\mathcal{M}_{k|k}^{x}$, $\mathcal{W}_{k}^{x}$ and $\mathcal{W}_{k}^{y}$, respectively, and $|A|$ is a matrix where each element is the absolute value of the corresponding element in $A$, i.e., $|A|=\left[|a_{i,j}|\right]$.
Proof 3.4.
@@
Since the volume of $\mathcal{Y}_{k|k}$ is the sum of $p_{k|k}^{y}$, $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ is also bounded by a function of the radius of the observation set $\mathcal{M}_{k|k}^{x}$. Consequently, if $\mathcal{M}_{k|k}^{x}$ is small, then the privacy level would be low, and the adversary retains little uncertainty after performing the inference attack.
@@
$\Delta\mathcal{Y}_{k|k}=\mathcal{Y}_{k|k-1}\setminus\mathcal{Y}_{k|k},$
is the difference between $\mathcal{Y}_{k|k-1}$ and $\mathcal{Y}_{k|k}$. Since $\mathcal{Y}_{k|k}$ is a subset of $\mathcal{Y}_{k|k-1}$, $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ can be computed with
$\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)=\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)-\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)\!.$ (42)
Therefore, to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$, it is equivalent to reduce the amount of uncertainty reduction $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ since the amount of prior uncertainty $\mathrm{Vol}\left(\mathcal{Y}_{k|k-1}\right)$ is fixed at time $k$. As shown in the next theorem, the amount of uncertainty reduction, i.e., $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$, is bounded by the new information extracted from $\mathcal{M}^{x}_{k|k}$.
@@
$\!\!\!\!\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)\!=\!\left\|\Psi\!\left(A_{3}\right)\!\Delta\mathcal{X}_{k-1|k}\oplus\Psi\!\left(A_{4}\right)\!\Delta\mathcal{Y}_{k-1|k}\right\|_{1}\!,\!\!$ (43)
@@
$\!\mathrm{Vol}\!\left(\Delta\mathcal{Y}_{k|k}\right)\!\!\leqslant$ $\!\left\|\!A_{3}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!+\!\left\|\!A_{4}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)\!,$
where $\Delta\mathcal{X}_{k-1|k}\!\!=\!\!\mathcal{X}_{k-1|k-1}\!\setminus\!\mathcal{X}_{k-1|k}$, $\Delta\!\mathcal{Y}_{k-1|k}\!\!=\!\!\mathcal{Y}_{k-1|k-1}\!\setminus\!\mathcal{Y}_{k-1|k}$ are the adversary’s uncertainty reduction of $X_{k-1}$ and $Y_{k-1}$, and $\left\|c_{k|k}^{y}\!-\!c_{k|k-1}^{y}\right\|_{1}$ quantifies the difference in central estimation with and without considering $\mathcal{X}_{k|k}$.
Proof 3.5.
@@
According to Theorem 9, the reduction of uncertainty $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k|k}\right)$ is highly correlated with the amount of information that the adversary extracts from the observation set $\mathcal{M}_{k|k}^{x}$. Furthermore, with (42) and Theorem 9, we can show the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k-1|k}\right)$ is bounded in the following lemma.
@@
$\mathrm{Vol}\!\left(\!\mathcal{Y}_{k|k-1}\!\right)\!-\!\left\|\!A_{3}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{X}_{k-1|k}\right)\!-\!\left\|\!A_{4}\!\right\|_{1}\!\mathrm{Vol}\!\left(\!\Delta\mathcal{Y}_{k-1|k}\right)$
@@
As a result, we can reduce the extracted information $\mathrm{Vol}\left(\Delta\mathcal{X}_{k-1|k}\right)$ and $\mathrm{Vol}\left(\Delta\mathcal{Y}_{k-1|k}\right)$ to increase the privacy level $\mathrm{Vol}\left(\mathcal{Y}_{k|k}\right)$ via designing proper observation set $\mathcal{M}^{x}_{k|k}$. Moreover, when the privacy level is high, the adversary’s ability to update its central estimate $c_{k|k}^{y}$ is also limited, as indicated by the small value of $\left\|c_{k|k}^{y}-c_{k|k-1}^{y}\right\|_{1}$. Based on this observation, $\mathcal{M}^{x}_{k|k}$ can also be designed to hinder accurate central estimate updates, further improving the privacy level.
@@ 4 Privacy Filter Design Problem Using the Volumetric Privacy measure
As discussed previously, an inappropriate choice of the observation set would cause privacy leakage of the private state through inference attacks. To mitigate this risk, we address the privacy filter design problem in this section. The proposed filter determines an appropriate observation set that achieves a desirable balance between preserving the data utility of the public state and ensuring the privacy protection of the private state.
@@ 4.1 The Structure of Privacy Filter
We begin by defining the decision domain of the privacy filter as follows. At time $k$, given the last decision set $\mathcal{X}_{k-1|k-1}$ and the private set $\mathcal{Y}_{k-1|k-1}$, the inference set of $X_{k}$ can be computed via,
$\mathcal{Y}_{k|k-1}=A_{3}\mathcal{X}_{k-1|k-1}\oplus A_{4}\mathcal{Y}_{k-1|k-1}\oplus B_{2}\mathcal{W}_{k}^{y},$
which contains all possible public states that can be reached from any states in $\mathcal{X}_{k-1|k-1}$ and $\mathcal{Y}_{k-1|k-1}$. Therefore, $\mathcal{X}_{k|k-1}$ is the maximum observation set $\mathcal{M}^{x}_{k|k}$ that the filter can release, i.e., $\mathcal{M}^{x}_{k|k}\subseteq\mathcal{X}_{k|k-1}$. To maintain the high data utility, the observation set has to satisfy the following constraint,
@@
To reduce privacy leakage while preserving data utility, we design the privacy filter illustrated in Fig. 2. The design consists of two steps: (1) randomly generate a set $\mathcal{S}_{k|k}^{x}$ such that $\mathcal{S}_{k|k}^{x}\subseteq\mathcal{X}_{k|k-1}$ and $\mathrm{Vol}(\mathcal{S}_{k|k}^{x})\leq\epsilon^{x}$; (2) optimize the observation set $\mathcal{M}^{x}_{k|k}$, which contains $\mathcal{S}_{k|k}^{x}$, to maximize the privacy level. Specifically, in the optimization step, given $\mathcal{S}_{k|k}^{x}$, we maximize the privacy level under the inference attack (6)-(12) by solving
$\mathbf{P_{1}}:$~ $\max_{\mathcal{M}^{x}_{k|k}}~\mathrm{Vol}\!\left(\mathcal{Y}_{k|k}\right)$ (44)
[…]
Note that $\mathcal{S}^{x}_{k|k}$ is randomly generated as a subset of $\mathcal{X}_{k|k-1}$. In practice, it can be sufficiently small; for instance, it may contain only the true public state $x_{k}$. Consequently, recovering $\mathcal{S}^{x}_{k|k}$ could lead to potential privacy leakage and should therefore be avoided. In the following, we show that an attacker cannot recover $\mathcal{S}^{x}_{k|k}$ by inverting the optimization problem $\mathbf{P_{1}}$ due to the randomization operation.
Proposition 11.
The attacker cannot obtain the smaller set $\mathcal{S}^{x}_{k|k}$ by inverting the optimization problem $\mathbf{P_{1}}$.
Proof 4.1.
First, $\mathcal{S}_{k|k}^{x}$ is selected as a random subset of $\mathcal{X}_{k|k-1}$ that contains the true state $x_{k}$. Consequently, $x_{k}$ may reside on the boundary of $\mathcal{S}_{k|k}^{x}$. Next, let $\mathcal{M}_{k|k}^{x,\star}$ denote the optimal observation set. In some cases, $\mathcal{S}_{k|k}^{x}$ may coincide with $\mathcal{M}_{k|k}^{x,\star}$, in which case $x_{k}$ may also lie on the boundary of $\mathcal{M}_{k|k}^{x,\star}$. Therefore, an attacker cannot reconstruct a strictly smaller feasible set containing $x_{k}$ by inverting the optimization process.
As a result, the proposed privacy filter exhibits the following properties: (1) In the absence of inference attacks, the filter output $\mathcal{M}^{x}_{k|k}$ satisfies the utility constraint. (2) In the presence of the inference attack described in (6)–(12), the filter output maximizes the privacy level. (3) The filter is robust against reverse attacks that attempt to recover the sufficiently small set $\mathcal{S}_{k|k}^{x}$.
Moreover, since the computational complexity of both the inference attack in (6)–(12) and the volume computation increases with the complexity of the set representations, a trade-off exists between the achievable privacy enhancement of the proposed filter and its computational cost. While more sophisticated set representations could improve the accuracy of privacy evaluation, for efficiency and clarity, we next present a concrete design based on the interval approximation described in Sec. 3.2.
Figure 2: Structure of the proposed privacy filter.